OpenStudy (anonymous):

Find the domain and range of the relation, and state whether or not the relation is a function. {(1, 3), (2, 3), (3, 3), (4, 3)}

OpenStudy (accessdenied):

The domain is the set of the first values in each ordered pair The range is the set of the second values in each ordered pair It would be a function if for each value in the domain, there is only one not necessarily unique corresponding range value.

OpenStudy (anonymous):

It is a function if every value of the domain has a unique value of the range. If x=y, then f(x) =f(y) for example if f(x) =2 and f(x) =5, then f is not a function. The relation in that question is a function. @AccessDenied the statement below is not true It would be a function if for each value in the domain, there is only one not necessarily unique corresponding range value.

OpenStudy (accessdenied):

What I meant was, for each x in the domain, there is a y-value that may or may not be the same as another x-value's, but you may not have two y-values for one x-value. I.e. f(x) = x^2, x=2 and x=-2 have the same y-value. Maybe my words are mixed up there, or I'm missing something...

OpenStudy (accessdenied):

As far as I can tell though, that definition is correct. o.O If each value \(x\) in the domain maps to \(\mathbf{one}\) (not necessarily unique) value \(y\) in the range, then the relation is a function. \( \begin{array}{|c|c|} \hline \mathbf x & \mathbf y \\ \hline 1 & 3\\ 2 & 3 \\ 3 & 3 \\ 4 & 3 \\ \hline \end{array} \qquad 1 \to 3,\ 2 \to 3,\ 3 \to 3,\ 4 \to 3\) The relation is a function because for each \(x\), there is one value of \(y\), which is not unique in this case between the other \(x\)'s.

OpenStudy (anonymous):

If a relation f is such f(1) =2 and f(1) =3 then f is not a function.

OpenStudy (accessdenied):

Yep, definitely true. Two y-values for one x-value doesn't fit that definition. I guess it may be because I say 'not necessarily unique" somewhat ambiguously to clarify that its the x-value that cannot have two y-values and not the y-value having two x-values that matters..